# Query Performance Prediction using Relevance Judgments Generated by Large Language Models

### v2

@@ 5.6. Implementation details
For all unsupervised QPP baselines, we tune the hyper-parameters for predicting the ranking quality of a ranker (either BM25 or ANCE) on TREC-DL 19 (TREC-DL 21) based on Pearson’s $\rho$ correlation for predicting the ranking quality of the same ranker on TREC-DL 20 (TREC-DL 22), and vice versa. We select the cut-off value $k$ for Clarity, NQC, WIG, SMV and so on from $\{5,10,15,20,25,50,100,300,500,1000\}$. n($\sigma_{x\%}$) has a a hyper-parameter $x$, which we choose from the set $\{0.25,0.4,0.5,0.6,0.75,0.9\}$.
To predict the performance of a certain ranker (either BM25 or ANCE), we train all supervised QPP baselines based on the ranked list returned by the target ranker. To predict a certain IR evaluation measure, regression-based methods (Khodabakhsh and Bagheri, 2023; Arabzadeh et al., 2021b; Hashemi et al., 2019) are trained to learn to output the target evaluation measure during training. However, our preliminary result shows that training supervised QPP baselines, especially for regression-based supervised methods (Khodabakhsh and Bagheri, 2023; Arabzadeh et al., 2021b; Hashemi et al., 2019), on the training set of MS MARCO V1 leads to inferior QPP quality for predicting the performance of ANCE. We hypothesize that this is because ANCE was originally trained on the training set of MS MARCO V1 (Xiong et al., 2021), and so the ranked list returned by ANCE on the training set of MS MARCO V1 would have higher quality than the ranked list returned by ANCE on the evaluation sets; therefore, supervised QPP methods that share the same training set as ANCE, tend to predict inflated performance on the evaluation sets, leading to degraded QPP quality. To solve the issue and ensure the consistency of the paper, we train all supervised QPP methods (including QPP-GenRE) on the development set of MS MARCO V1 (6980 queries) for predicting the performance of either BM25 or ANCE. We train all supervised QPP methods for 5 epochs and pick the best checkpoint for predicting the performance of a ranker on TREC-DL 19 (TREC-DL 21) based on Pearson’s $\rho$ correlation for predicting the performance of the same ranker on TREC-DL 20 (TREC-DL 22) and vice versa. All supervised QPP baselines use bert-base-uncased, [^5] https://github.com/huggingface/transformers a constant learning rate (0.00002), and the Adam optimizer (Kingma and Ba, 2015).
@@
We equip QPP-GenRE with LLaMA-7B for judging relevance. We use a recent PEFT method, 4-bit QLoRA (Dettmers et al., 2023), to fine-tune LLaMA-7B. The training of judging relevance needs positive and negative items per query. For positive items, we use the items annotated as relevant in qrels per query; we randomly sample one negative item from the ranked list (1,000 items) returned by BM25 per query. We fine-tune LLaMA-7B for 5 epochs on the development set of MS MARCO V1, taking about an hour and a half on an NVIDIA A100 GPU (40GB).
@@
Table 2. Correlation coefficients (Pearson’s $\rho$ and Kendall’s $\tau$) between actual retrieval quality, in terms of RR@10, of ANCE and performance predicted by QPP-GenRE/baselines, on TREC-DL 19 and 20. indicates statistically significant correlation coefficients ($p$-value $<0.05$). indicates the statistically significant improvement of QPP-GenRE compared to all the baselines (paired $t$-test; $p$-value $<0.001$ with Bonferroni correction for multiple testing). The best value in each column is marked in bold. $n$ denotes QPP-GenRE’s judgment depth in a ranked list.
QPP method Ranker: ANCE
TREC-DL 19 TREC-DL 20
P-$\rho$ K-$\tau$ P-$\rho$ K-$\tau$
Clarity -0.078 -0.012 -0.074 -0.048
WIG 0.313 0.228 0.059 0.048
NQC 0.350 0.200 0.145 0.112
$\sigma_{max}$ 0.384 0.287 0.171 0.118
n($\sigma_{x\%}$) 0.200 0.176 -0.008 0.022
SMV 0.352 0.256 0.182 0.161
UEF(NQC) 0.340 0.260 0.131 0.108
RLS(NQC) 0.359 0.273 0.178 0.139
QPP-PRP 0.259 0.246 0.100 -0.008
Dense-QPP 0.452 0.280 0.209 0.139
NQA-QPP -0.026 -0.009 -0.059 -0.080
BERTQPP 0.330 0.214 0.046 -0.012
qppBERT-PL 0.092 0.025 -0.224 -0.218
M-QPPF 0.292 0.200 0.068 0.038
QPP-LLM (few-shot) -0.008 0.005 -0.226 -0.207
QPP-LLM (fine-tuned) -0.073 0.011 -0.022 0.069
QPP-GenRE ($n=10$) 0.567 0.440 0.293 0.257
@@ 6.1. Predicting a precision-oriented IR measure
To answer RQ1, we compare QPP-GenRE and all baselines in predicting the performance of BM25 and ANCE w.r.t. a widely-used precision-oriented metric, RR@10; see Table 2. We have three main observations.
First, our proposed method, QPP-GenRE, outperforms all baselines in terms of both correlation coefficients on all datasets when predicting the performance of both rankers. In particular, we observe that QPP-GenRE outperforms QPP-PRP (Singh et al., 2023), which is a recently proposed baseline by 84% (0.292 vs. 0.538) in terms of Pearson’s $\rho$ when predicting RR@10 for BM25 on TREC-DL 19.
@@
QPP method TREC-DL 19 TREC-DL 20 TREC-DL 21 TREC-DL 22
P-$\rho$ K-$\tau$ P-$\rho$ K-$\tau$ P-$\rho$ K-$\tau$ P-$\rho$ K-$\tau$
@@
Table 4. Correlation coefficients (Pearson’s $\rho$ and Kendall’s $\tau$) between actual retrieval quality, in terms of nDCG@10, of ANCE and performance predicted by QPP-GenRE/baselines, on TREC-DL 19 and 20. $n$ denotes QPP-GenRE’s judgment depth in a ranked list. indicates statistically significant correlation coefficients ($p$-value $<0.05$). indicates the statistically significant improvement of QPP-GenRE ($n$=200) compared to all the baselines (paired $t$-test; $p$-value $<0.001$ with Bonferroni correction for multiple testing). The best value in each column is marked in bold.
QPP method TREC-DL 19 TREC-DL 20
P-$\rho$ K-$\tau$ P-$\rho$ K-$\tau$
Clarity -0.088 -0.062 -0.091 -0.045
WIG 0.515 0.368 0.218 0.150
NQC 0.548 0.372 0.411 0.290
$\sigma_{max}$ 0.455 0.339 0.403 0.288
n($\sigma_{x\%}$) 0.388 0.315 0.103 0.075
SMV 0.496 0.359 0.380 0.283
UEF(NQC) 0.548 0.372 0.413 0.290
RLS(NQC) 0.466 0.346 0.333 0.271
QPP-PRP 0.129 0.049 0.216 0.121
Dense-QPP 0.565 0.389 0.419 0.318
NQA-QPP 0.089 -0.038 0.186 0.113
BERTQPP 0.222 0.117 0.137 0.089
qppBERT-PL 0.116 0.098 -0.119 -0.046
M-QPPF 0.287 0.160 0.225 0.177
QPP-LLM (few-shot) 0.136 0.120 -0.130 -0.094
QPP-LLM (fine-tuned) 0.203 0.117 0.081 0.097
QPP-GenRE ($n=200$) 0.712 0.483 0.457 0.343
QPP-GenRE ($n=10$) 0.624 0.406 0.306 0.238
QPP-GenRE ($n=100$) 0.719 0.489 0.456 0.355
QPP-GenRE ($n=1,000$) 0.719 0.492 0.447 0.321
@@ 6.2. Predicting an IR measure considering recall
To answer RQ2, Table 4 lists the performance of QPP-GenRE along with all the baselines on assessing BM25 and ANCE in terms of nDCG@10. For QPP-GenRE, we universally set the judging depth $n$ to 200 for all evaluation sets. The result reveals that by judging only 200 items per query, we can achieve state-of-the-art QPP quality in terms of nDCG@10 for all rankers on all evaluation sets; we will investigate the impact of judging depth on QPP-GenRE’s performance in the next section. Also, QPP-LLM (few-shot) and QPP-LLM (fine-tuning) are among the worst-performing baselines, showing that the LLMs struggle to generate numerical scores. Different from the results for RQ1, most QPP methods tend to perform better when predicting nDCG@10 than RR@10; this observation indicates that predicting RR@10 is a more challenging task.
@@ 7.1. Judging depth analysis
To answer RQ3, as detailed in Section 4.3.2, for predicting IDCG, we devise an approximation strategy and use the items in the top $n$ ranks of the ranked list $L$ that are predicted as relevant by QPP-GenRE to approximate all the relevant items for a query in the corpus. To investigate the impact of the value of $n$ on the quality of the prediction, we investigate the relationship between the QPP quality of predicting nDCG@10 and the judgment depth to answer the following question: What depth of relevance judgment $n$ do we need to consider to get a satisfactory performance for predicting nDCG@10? In Figure 3, we plot the correlation coefficients between actual nDCG@10 values and nDCG@10 values predicted by QPP-GenRE for different judging depths in {10, 25, 50, 75, 100, 200, 300, 400, 500, 600, 700, 800, 900, 1,000} on TREC-DL 19 and 20. We also show exact QPP results with depths at 10, 100 and 1000 in Table 4.
Table 4 reveals that, by judging only 10 items in the ranked list of BM25 and ANCE, we can already outperform all the baselines and achieve state-of-the-art QPP quality on half of the evaluation sets we used, e.g., assessing BM25 on TREC-DL 19 and 21, and assessing ANCE on TREC-DL 19. While judging deeper in the ranked list is essential for predicting recall-oriented measures, satisfactory QPP quality is still attainable with a relatively shallow depth. Moreover, Figure 3 illustrates that judging the top 200 items in a ranked list already reaches the saturation point for assessing BM25, i.e., there is no significant improvement by judging a higher number of items, while judging less than 100 top items reaches the saturation point for ANCE. We speculate that this is because ANCE has better retrieval quality than BM25, and more relevant items would appear earlier in the ranked list of ANCE than BM25; therefore, a shallower judging depth suffices to approximate all relevant items in the corpus. This emphasizes the need to consider retrieval quality when determining the optimal judgment depth for various rankers.
Table 5. Relevance judgment agreement (Cohen’s $\kappa$) between TREC assessors and each LLM, and Pearson’s $\rho$ correlation coefficients between BM25’ actual nDCG@10 values and those predicted by QPP-GenRE integrated with each LLM on TREC-DL 19–22. The best value in each column is marked in bold.
LLM TREC-DL 19 TREC-DL 20 TREC-DL 21 TREC-DL 22
$\kappa$ P-$\rho$ $\kappa$ P-$\rho$ $\kappa$ P-$\rho$ $\kappa$ P-$\rho$

### v3

@@ 5.6. Implementation details
For all unsupervised QPP baselines, we tune the hyper-parameters for predicting the ranking quality of a ranker (either BM25 or ANCE) on TREC-DL 19 (TREC-DL 21) based on Pearson’s $\rho$ correlation for predicting the ranking quality of the same ranker on TREC-DL 20 (TREC-DL 22), and vice versa. We select the cut-off value $k$ for Clarity, NQC, WIG, SMV and so on from $\{5,10,15,20,25,50,100,300,500,1000\}$. n($\sigma_{x\%}$) has a hyper-parameter $x$, which we choose from the set $\{0.25,0.4,0.5,0.6,0.75,0.9\}$.
To predict the performance of a certain ranker (any of BM25, ANCE, or TAS-B), we train all supervised QPP baselines based on the ranked list returned by the target ranker. To predict a certain IR evaluation measure, regression-based methods (Khodabakhsh and Bagheri, 2023; Arabzadeh et al., 2021b; Hashemi et al., 2019) are trained to learn to output the target evaluation measure during training. However, our preliminary result shows that training supervised QPP baselines, especially for regression-based supervised methods (Khodabakhsh and Bagheri, 2023; Arabzadeh et al., 2021b; Hashemi et al., 2019), on the training set of MS MARCO V1 leads to inferior QPP quality for predicting the performance of the neural rankers (ANCE and TAS-B). We hypothesize that this is because they were originally trained on the training set of MS MARCO V1 (Xiong et al., 2021; Hofstätter et al., 2021), and so the ranked list returned by them on the training set of MS MARCO V1 would have higher quality than the ranked list returned by them on the evaluation sets; therefore, supervised QPP methods that share the same training set as the neural rankers, tend to predict inflated performance on the evaluation sets, leading to degraded QPP quality. To solve the issue and ensure the consistency of the paper, we train all supervised QPP methods (including QPP-GenRE) on the development set of MS MARCO V1 (6980 queries) for predicting the performance of BM25, ANCE or TAS-B. We train all supervised QPP methods for 5 epochs and pick the best checkpoint for predicting the performance of a ranker on TREC-DL 19 (TREC-DL 21) based on Pearson’s $\rho$ correlation for predicting the performance of the same ranker on TREC-DL 20 (TREC-DL 22) and vice versa. All supervised QPP baselines use bert-base-uncased, [^4] https://github.com/huggingface/transformers a constant learning rate (0.00002), and the Adam optimizer (Kingma and Ba, 2015).
@@
We equip QPP-GenRE with an LLM for judging relevance. We use a recent PEFT method, 4-bit QLoRA (Dettmers et al., 2023), to fine-tune an LLM. To maintain a comparable setup with the baselines, we fine-tune an LLM for 5 epochs on the development set of MS MARCO V1. Note that we use LLaMA-7B for BM25 and ANCE, and Mistral-7B-Instruct-v0.3 for TAS-B. The training of judging relevance needs positive and negative items per query. For positive items, we use the items annotated as relevant in qrels per query; we randomly sample one negative item from the ranked list (1,000 items) returned by BM25 per query. There are 6,980 queries in our training set. Each each query may have multiple relevant items annotated in the qrels, and has one negative item we sampled. As a result, we have 7,437 positive training examples and 6,980 negative training examples. All experiments are conducted on an NVIDIA A100 GPU (40GB).
One might argue why we choose QLoRA fine-tuning instead of distilling an oracle model into a smaller model. (Sun et al., 2023b). The decision is based on the following three reasons. First, model distillation requires an existing oracle model, such as GPT-4, for relevance prediction. However, this work focuses on exclusively using open-source LLMs, avoiding using powerful commercial models like GPT-4 to ensure reproducibility and deterministic outputs. Second, one might wonder why we did not distill larger open-source LLMs into smaller models. As demonstrated in Figure 5, a 1-billion-parameter Llama model fine-tuned using QLoRA on human-labeled relevance judgments outperforms a 70-billion-parameter Llama model using few-shot prompting. Therefore, if we choose to distill the 70-billion-parameter model into a smaller model, the performance of the distilled model would be inferior to that achieved by the 1-billion-parameter model fine-tuned via QLoRA, because the performance of distilled model is inherently limited by the larger model’s capabilities. Third, we have a large amount of human-labeled relevance judgments available. Directly using these labels to fine-tune LLMs via QLoRA allows us to make the most efficient use of this data.
@@
Table 2. Correlation coefficients (Pearson’s $\rho$ and Kendall’s $\tau$) between actual retrieval quality, in terms of RR@10, of ANCE/TAS-B and performance predicted by QPP-GenRE/baselines, on TREC-DL 19 and 20. indicates statistically significant correlation coefficients ($p$-value $<0.05$). indicates the statistically significant improvement of QPP-GenRE compared to all the baselines (paired $t$-test; $p$-value $<0.001$ with Bonferroni correction for multiple testing). The best value in each column is marked in bold. $n$ denotes QPP-GenRE’s judgment depth in a ranked list.
Ranker: ANCE Ranker: TAS-B
TREC-DL 19 TREC-DL 20 TREC-DL 19 TREC-DL 20
QPP method P-$\rho$ K-$\tau$ P-$\rho$ K-$\tau$ P-$\rho$ K-$\tau$ P-$\rho$ K-$\tau$
Clarity -0.078 -0.012 -0.074 -0.048 -0.212 -0.148 0.148 0.133
WIG 0.313 0.228 0.059 0.048 -0.066 -0.125 0.024 0.020
NQC 0.350 0.200 0.145 0.112 0.248 0.213 0.260 0.194
$\sigma_{max}$ 0.384 0.287 0.171 0.118 0.015 0.021 0.312 0.245
n($\sigma_{x\%}$) 0.200 0.176 -0.008 0.022 -0.030 -0.079 0.080 0.086
SMV 0.352 0.256 0.182 0.161 0.249 0.205 0.263 0.198
UEF(NQC) 0.340 0.260 0.131 0.108 0.260 0.228 0.281 0.213
RLS(NQC) 0.359 0.273 0.178 0.139 0.257 0.217 0.283 0.217
QPP-PRP 0.259 0.246 0.100 -0.008 0.155 0.113 0.203 0.116
Dense-QPP 0.452 0.280 0.209 0.139 0.251 0.213 0.146 0.012
NQA-QPP -0.026 -0.009 -0.059 -0.080 0.172 0.144 -0.058 -0.075
BERTQPP 0.330 0.214 0.046 -0.012 0.202 0.194 0.077 0.037
qppBERT-PL 0.092 0.025 -0.224 -0.218 0.276 0.269 0.004 -0.002
M-QPPF 0.292 0.200 0.068 0.038 0.277 0.236 0.103 0.022
QPP-LLM (few-shot) -0.008 0.005 -0.226 -0.207 -0.080 0.002 0.054 -0.024
QPP-LLM (fine-tuned) -0.073 0.011 -0.022 0.069 0.155 0.113 0.043 -0.020
QPP-GenRE ($n=10$) 0.567 0.440 0.293 0.257 0.538 0.481 0.356 0.289
@@ 6.1. Predicting a precision-oriented IR measure
To answer RQ1, we compare QPP-GenRE and all baselines in predicting the performance of BM25, ANCE and TAS-B w.r.t. a widely-used precision-oriented metric, RR@10; see Tables 1 and 2. We have three main observations.
First, our proposed method, QPP-GenRE, outperforms all baselines in terms of both correlation coefficients on all datasets when predicting the performance of all rankers. In particular, we observe that QPP-GenRE outperforms QPP-PRP (Singh et al., 2023), which is a recently proposed baseline by 84% (0.292 vs. 0.538) in terms of Pearson’s $\rho$ when predicting RR@10 for BM25 on TREC-DL 19.
@@
Ranker: BM25
TREC-DL 19 TREC-DL 20 TREC-DL 21 TREC-DL 22
QPP method P-$\rho$ K-$\tau$ P-$\rho$ K-$\tau$ P-$\rho$ K-$\tau$ P-$\rho$ K-$\tau$
@@
Table 4. Correlation coefficients (Pearson’s $\rho$ and Kendall’s $\tau$) between actual retrieval quality, in terms of nDCG@10, of ANCE/TAS-B and performance predicted by QPP-GenRE/baselines, on TREC-DL 19 and 20. $n$ denotes QPP-GenRE’s judgment depth in a ranked list. indicates statistically significant correlation coefficients ($p$-value $<0.05$). indicates the statistically significant improvement of QPP-GenRE ($n$=200) compared to all the baselines (paired $t$-test; $p$-value $<0.001$ with Bonferroni correction for multiple testing). The best value in each column is marked in bold.
Ranker: ANCE Ranker: TAS-B
TREC-DL 19 TREC-DL 20 TREC-DL 19 TREC-DL 20
QPP method P-$\rho$ K-$\tau$ P-$\rho$ K-$\tau$ P-$\rho$ K-$\tau$ P-$\rho$ K-$\tau$
Clarity -0.088 -0.062 -0.091 -0.045 0.153 0.049 0.162 0.087
WIG 0.515 0.368 0.218 0.150 0.228 0.146 0.227 0.169
NQC 0.548 0.372 0.411 0.290 0.330 0.233 0.406 0.264
$\sigma_{max}$ 0.455 0.339 0.403 0.288 0.220 0.126 0.428 0.284
n($\sigma_{x\%}$) 0.388 0.315 0.103 0.075 -0.008 -0.031 0.002 -0.020
SMV 0.496 0.359 0.380 0.283 0.349 0.253 0.425 0.285
UEF(NQC) 0.548 0.372 0.413 0.290 0.321 0.246 0.425 0.271
RLS(NQC) 0.466 0.346 0.333 0.271 0.314 0.246 0.404 0.272
QPP-PRP 0.129 0.049 0.216 0.121 0.220 0.126 0.267 0.237
Dense-QPP 0.565 0.389 0.419 0.318 0.429 0.244 0.126 0.012
NQA-QPP 0.089 -0.038 0.186 0.113 -0.020 0.060 0.031 0.024
BERTQPP 0.222 0.117 0.137 0.089 0.043 0.027 0.178 0.086
qppBERT-PL 0.116 0.098 -0.119 -0.046 0.304 0.187 0.057 0.057
M-QPPF 0.287 0.160 0.225 0.177 0.163 0.051 0.304 0.171
QPP-LLM (few-shot) 0.136 0.120 -0.130 -0.094 -0.020 0.060 0.108 0.048
QPP-LLM (fine-tuned) 0.203 0.117 0.081 0.097 0.262 0.195 0.162 0.111
QPP-GenRE ($n=200$) 0.712 0.483 0.457 0.343 0.501 0.346 0.449 0.315
QPP-GenRE ($n=10$) 0.624 0.406 0.306 0.238 0.490 0.309 0.421 0.290
QPP-GenRE ($n=100$) 0.719 0.489 0.456 0.355 0.501 0.336 0.450 0.317
QPP-GenRE ($n=1,000$) 0.719 0.492 0.447 0.321 0.505 0.348 0.449 0.315
@@ 6.2. Predicting an IR measure considering recall
To answer RQ2, Tables 3 and 4 list the performance of QPP-GenRE along with all the baselines on assessing BM25, ANCE and TAS-B in terms of nDCG@10. For QPP-GenRE, we universally set the judging depth $n$ to 200 for all evaluation sets. The result reveals that by judging only 200 items per query, we can achieve state-of-the-art QPP quality in terms of nDCG@10 for all rankers on all evaluation sets; we will investigate the impact of judging depth on QPP-GenRE’s performance in the next section. Also, QPP-LLM (few-shot) and QPP-LLM (fine-tuning) are among the worst-performing baselines, showing that the LLMs struggle to generate numerical scores. Different from the results for RQ1, most QPP methods tend to perform better when predicting nDCG@10 than RR@10; this observation indicates that predicting RR@10 is a more challenging task.
@@ 7.1. Judging depth analysis
RQ3 examines how varying the number of top-ranked documents submitted for relevance judgments impacts QPP quality. To answer RQ3, as detailed in Section 4.3.2, for predicting IDCG, we devise an approximation strategy and use the items in the top $n$ ranks of the ranked list $L$ that are predicted as relevant by QPP-GenRE to approximate all the relevant items for a query in the corpus. To investigate the impact of the value of $n$ on the quality of the prediction, we investigate the relationship between the QPP quality of predicting nDCG@10 and the judgment depth to answer the following question: What depth of relevance judgment $n$ do we need to consider to get a satisfactory performance for predicting nDCG@10? In Figure 4, we plot the correlation coefficients between actual nDCG@10 values and nDCG@10 values predicted by QPP-GenRE for different judging depths in {10, 25, 50, 75, 100, 200, 300, 400, 500, 600, 700, 800, 900, 1,000} on TREC-DL 19 and 20. We also show exact QPP results with depths at 10, 100 and 1000 in Tables 3 and 4.
Tables 3 and 4 reveal that, by judging only 10 items in the ranked list, we can already outperform all the baselines and achieve state-of-the-art QPP quality on half of the evaluation sets we used, e.g., assessing BM25 on TREC-DL 19/21, ANCE on TREC-DL 19, and TAS-B on TREC-DL 19. While judging deeper in the ranked list is essential for predicting recall-oriented measures, satisfactory QPP quality is still attainable with a relatively shallow depth. Moreover, Figure 4 illustrates that judging the top 200 items in a ranked list already reaches the saturation point for assessing BM25, i.e., there is no significant improvement by judging a higher number of items, while judging less than 100 top items reaches the saturation point for ANCE. We speculate that this is because ANCE has better retrieval quality than BM25, and more relevant items would appear earlier in the ranked list of ANCE than BM25; therefore, a shallower judging depth suffices to approximate all relevant items in the corpus. This emphasizes the need to consider retrieval quality when determining the optimal judgment depth for various rankers.
Table 5. Relevance judgment agreement (Cohen’s $\kappa$) between TREC assessors and each LLM, and Pearson’s $\rho$ correlation coefficients between BM25’s actual nDCG@10 values and those predicted by QPP-GenRE integrated with each LLM on TREC-DL 19–22. The best value in each column is marked in bold. We do not fine-tune Llama-3-70B-Instruct due to budget constraints.
TREC-DL 19 TREC-DL 20 TREC-DL 21 TREC-DL 22
LLM $\kappa$ P-$\rho$ $\kappa$ P-$\rho$ $\kappa$ P-$\rho$ $\kappa$ P-$\rho$
@@
Llama-3.2-1B-Instruct (few-shot) 0.013 0.152 0.029 0.099 0.009 0.249 0.079 0.087
Llama-3.2-3B-Instruct (few-shot) 0.186 0.293 0.114 0.020 0.165 0.289 0.055 0.443
Mistral-7B-Instruct-v0.3 (few-shot) 0.224 0.271 0.174 0.499 0.245 0.414 0.042 0.243
@@
Mistral-22B-Instruct (few-shot) 0.281 0.412 0.238 0.535 0.276 0.528 0.083 0.473
Llama-3-70B-Instruct (few-shot) 0.321 0.526 0.245 0.557 0.279 0.545 0.086 0.483
Llama-3.2-1B-Instruct (fine-tuned) 0.351 0.610 0.211 0.596 0.197 0.570 0.042 0.428
Llama-3.2-3B-Instruct (fine-tuned) 0.383 0.710 0.273 0.722 0.306 0.608 0.042 0.511
Mistral-7B-Instruct-v0.3 (fine-tuned) 0.403 0.734 0.328 0.720 0.373 0.592 0.076 0.411
@@
Mistral-22B-Instruct (fine-tuned) 0.407 0.682 0.276 0.640 0.310 0.591 0.071 0.462
(a) TREC-DL 19
(b) TREC-DL 20
(c) TREC-DL 21
(d) TREC-DL 22
Figure 5. Pearson’s $\rho$ correlation coefficients between BM25’ actual nDCG@10 values and those predicted by QPP-GenRE integrated with various LLMs with different sizes, under both few-shot and fine-tuned settings, on TREC-DL 19–22. From left to right along the x-axis, as the size of the LLMs increases, the inference efficiency correspondingly decreases. To aid visual comparison, LLMs from the same family share a consistent colour scheme: Llama models are shown in blue, and Mistral models in orange. Note that, for simplicity, we only retain the results of LLMs with “instruct” versions. We do not fine-tune Llama-3-70B-Instruct due to budget constraints.